# Data-Driven LQR with Finite-Time Experiments via Extremum-Seeking Policy Iteration

### v2

@@ V-B Averaged System Analysis
$\sum_{\tau=k+1}^{k+\mathbf{k}_{\text{prd}}}(J_{T}(K+\delta\mathrm{D}^{\tau})-\mathrm{z})$
@@
$\leq\exp(-\gamma ak)V_{\lambda}(\mathrm{z}_{\textsc{av}}^{0}-J_{\textsc{av}}^{\delta}(K_{\textsc{av}}^{0}),K_{\textsc{av}}^{0}),$ (19)
for all $V_{\lambda}(\mathrm{z}_{\textsc{av}}^{k}-J_{\textsc{av}}^{\delta}(K_{\textsc{av}}^{k}),K_{\textsc{av}}^{k})\geq\rho_{\textsc{av}}$. $\blacksquare$
@@ V-C Proof of Theorem 2
The proof relies on the application of Theorem 1 (cf. Section II) to system (10). Then, in order to apply Theorem 1, we need to (i) choose the design parameters $c_{1},\rho>0$ bounding the initial and final values of $V_{\lambda}$, respectively, and (ii) satisfy the conditions required by Assumptions 12, and 3. By [50, Lemma 3.8], we recall that there exists $\psi>0$ such that
$\psi\left\|K-K^{\star}\right\|\leq J(K)-J(K^{\star}),$ (20)
for all $K\in\mathcal{K}$. Therefore, by looking at the statement of Theorem 1 and given the desired final radius $r$, we set $\rho\in(0,r\psi]$. In order to set the initial radius, we need to find a bound for $\delta$ such that $K^{0}+\delta\mathrm{D}^{k}$ is stabilizing for all $k\in\mathbb{N}$. To this end, we note that $K^{0}\in\mathcal{K}$, $\mathcal{K}$ is open [49, Lemma IV.3], and $\mathrm{D}^{k}$ is bounded for all $k\in\mathbb{N}$. Hence, there exists $\bar{\delta}_{0}>0$ such that $K^{0}+\delta\mathrm{D}^{k}\in\mathcal{K}$ for all $\delta\in[0,\bar{\delta}]$ and $k\in\{1,\ldots,\mathbf{k}_{\text{prd}}\}$. Now, we arbitrarily choose $c_{1}>\max_{\delta\in[0,\bar{\delta}_{0}]}V_{1}(\mathrm{z}^{0}-J_{\textsc{av}}^{\delta}(K^{0}),K^{0})$ and, thus, we note that $c_{1}\geq\max_{\delta\in[0,\bar{\delta}]}V_{\lambda}(\mathrm{z}^{0}-J_{\textsc{av}}^{\delta}(K^{0}),K^{0})$ for all $\lambda\geq 1$ (see the definition of $V_{\lambda}$ in (18)). Once the initial and final radius $c_{1}$ and $\rho$ have been chosen, let us check Assumptions 12, and 3. First, Assumption 1 is trivially satisfied because the dither signals are $\mathbf{k}_{\text{prd}}$-periodic (cf. Assumption 6). Second, we remark that (10) and its corresponding averaged system (16) are continuous over the set $\{(\mathrm{z},K)\in\mathbb{R}\times\mathcal{K}\mid K+\delta\mathrm{D}^{k}\in\mathcal{K}\text{ for all }k\in\mathbb{N}\}$, as required by Assumption 2. For this reason, let us choose $\delta$ such that the level set $\Omega_{c_{1}}:=\{(\mathrm{z},K)\in\mathbb{R}\times\mathcal{K}\mid V_{1}(\mathrm{z}-J_{\textsc{av}}^{\delta}(K),K)\leq c_{1}\}$ of $V_{1}$ (i.e, the function $V_{\lambda}$ with $\lambda=1$, see (18)) is contained into $\{(\mathrm{z},K)\in\mathbb{R}\times\mathcal{K}\mid K+\delta\mathrm{D}^{k}\in\mathcal{K}\text{ for all }k\in\mathbb{N}\}$. To this end, by looking at the definition of $V_{\lambda}$ (cf. (18)), we note that
$V_{1}(\mathrm{z}-J_{\textsc{av}}^{\delta}(K),K)\leq c_{1}\implies J(K)-J(K^{\star})\leq c_{1},$ (21)
@@
$(\mathrm{z},K)\in\Omega_{c_{1}}\implies K\in\mathcal{K}.$
Moreover, we recall that $\mathcal{K}$ is open [49, Lemma IV.3] and $\mathrm{D}^{k}$ is bounded for all $k\in\mathbb{N}$. Then, we guarantee the existence of $\bar{\delta}_{2}>0$ such that, for all $\delta\in(0,\min\{\bar{\delta}_{0},\bar{\delta}_{2}\})$, it holds $K+\delta\mathrm{D}^{k}\in\mathcal{K}$ for all $K$ satisfying $J(K)-J(K^{\star})\leq c_{1}$ and $k\in\mathbb{N}$. With these results at hand, Lemma 3 ensures the existence of $\bar{\gamma}_{1},\bar{\delta}_{1}>0$, $\bar{\lambda}\geq 1$, and $\bar{T}\in\mathbb{N}$ such that, by setting $\gamma\in(0,\bar{\gamma}_{1})$, $\delta\in(0,\bar{\delta})$ with $\bar{\delta}:=\min\{\bar{\delta}_{0},\bar{\delta}_{1},\bar{\delta}_{2}\}$, $T>\bar{T}$ and $\lambda\geq\bar{\lambda}$, $V_{\lambda}$ achieves the convergence properties (19) along the trajectories of the averaged system (16) and, thus, Assumption 3 is satisfied. Hence, we are entitled to apply Theorem 1. Specifically, it ensures the existence of $\bar{\gamma},\bar{k}>0$ such that, for all $\gamma\in(0,\bar{\gamma})$ and $k\geq\bar{k}$, it holds
$V_{\lambda}(\mathrm{z}^{k}-J(K^{k}),K^{k})\leq\rho,$ (22)
for all $(\mathrm{z}^{0},K^{0})\in\Omega_{c_{1}}$. The proof of (12) follows by combining $J(K)-J(K^{\star})\leq V(\mathrm{z}-J(K),K)$, (20), (22), and the choice of $\rho$.
@@ VI Numerical Simulations: Control of a Doubly Fed Induction Motor
In this section, we numerically test the effectiveness of EXP-LQR. To this end, we consider a forward Euler discretization of the continuous-time linear model provided by [51] for a Doubly Fed Induction Motor (DFIM) operating at constant speed. Namely, we consider the discrete-time linear system
@@
where $\Delta t=10^{-2}$ is the adopted sampling period, while $x,u\in\mathbb{R}^{4}$ are the state and input variables, respectively, and are defined as
$x:=\begin{bmatrix}i_{1u}\\ i_{1v}\\ i_{2u}\\ i_{2v}\end{bmatrix},\quad u:=\begin{bmatrix}u_{1u}\\ u_{1v}\\ u_{2u}\\ u_{2v}\end{bmatrix},$
@@
$L_{1}$ [htpb] $0.02645$ $R_{1}$ [$\Omega$] $0.036$
$L_{2}$ [htpb] $0.0264$ $R_{2}$ [$\Omega$] $0.038$
$L_{m}$ [htpb] $0.0257$ $\omega_{0}$ [rad/s] $2\pi 70.8$
@@
For the cost matrices $Q\in\mathbb{R}^{4\times 4}$ and $R\in\mathbb{R}^{4\times 4}$, we randomly generate them to ensure they are symmetric, with eigenvalues lying within the interval $(0,2)$ thus satisfying Assumption 4.
We arbitrarily pick $T=20$, i.e., the algorithm needs to perform $4$ experiments or simulations given by $20$ samples per iteration to retrieve the truncated cost $J_{T}(K^{k}+\delta\mathrm{D}^{k})$, see Remark 1. We empirically tune the other algorithm parameters as $\gamma=10^{-8}$ and $\delta=10^{-2}$. As for the generation of the dither matrix $\mathrm{D}^{k}\in\mathbb{R}^{4\times 4}$, we ordered the pairs $(i,j)\in\{1,\ldots,4\}\times\{1,\ldots,4\}$ with indices $p=1,\dots,16$ and chosen $\mathbf{k}_{p,\text{prd}}=19^{-(p-1)/2}$ and $\phi_{p}=0$ for $p$ odd, while $\mathbf{k}_{p,\text{prd}}=\mathbf{k}_{p-1,\text{prd}}$ and $\phi_{p}=\pi/2$ for $p$ even. This choice ensures that Assumption 6 is satisfied with period $\mathbf{k}_{\text{prd}}=19$.
Fig. 4 shows the evolution of the relative cost error $(J(K^{k})-J(K^{\star}))/J(K^{\star})$ along the algorithm iterations $k$ in logarithmic scale. As predicted by Theorem 2, Fig. 4 shows that EXP-LQR asymptotically converges in a neighborhood of the optimal gain $K^{\star}$.
Figure 4: Evolution of the relative cost error $(J(K^{k})-J(K^{\star}))/J(K^{\star})$ along the algorithm iteration $k$.
In Fig. 5, we show the evolution of $\sigma_{\text{max}}(A+B(K^{k}+\delta\mathrm{D}^{k}))$ along the algorithm iterations $k$, where, given a generic square matrix $M\in\mathbb{R}^{n\times n}$, the symbol $\sigma_{\text{max}}(M)$ denotes the maximum (in absolute value) eigenvalue of $M$. In particular, Fig. 5 shows that $\sigma_{\text{max}}(A+B(K^{k}+\delta\mathrm{D}^{k}))$ never reaches the unitary value, i.e., we always test the system through a stabilizing state-feedback controller $u_{t}=(K^{k}+\delta\mathrm{D}^{k})x_{t}$.
@@
Finally, in Fig. 6, we show the evolution of the norm of the state trajectory $\left\|x_{t}\right\|$ of system (23) in four simulations (each composed of $T=20$ samples) performed at different algorithm iterations $k$ to retrieve the truncated cost $J_{T}(K^{k}+\delta\mathrm{D}^{k})$. In particular, Fig. 6 shows that the trajectories of system (23) (controlled with $u_{t}=(K^{k}+\delta\mathrm{D}^{k})u_{t}$) exponentially converge to the origin quicker and quicker as the iteration index $k$ increases since we are iteratively reducing the absolute values of the eigenvalues of the gain closed-loop matrix $(A+BK^{k})$ (see also Fig. 5).
@@ VII Conclusions
We proposed EXP-LQR, i.e., a novel data-driven method able to iteratively find the state feedback gain matrix solving a Linear Quadratic Regulator problem. EXP-LQR does not need the direct knowledge of the system matrices, cost matrices, and state measurements. Indeed, given an oracle able to provide a finite-time truncation of the LQR cost, our method refines its estimate according to a mechanism based on extremum-seeking. We analyzed the resulting time-varying algorithm by exploiting system theory tools based on Lyapunov stability and averaging theory. Specifically, we guaranteed that our algorithm exponentially converges to an arbitrarily small ball containing the optimal gain matrix. We tested the proposed solution with numerical simulations involving the control of an inductance motor.
@@
$\left\|f(\chi,k)\right\|$ $\leq L$ (26a)
$\left\|f_{\textsc{av}}(\chi)\right\|$ $\leq L$ (26b)
$\left\|\dfrac{\partial f(\chi,k)}{\partial\chi}\right\|$ $\leq L$ (26c)
$\left\|\nabla f_{\textsc{av}}(\chi)\right\|$ $\leq L$ (26d)
$\left\|\nabla V(\chi)\right\|$ $\leq L,$ (26e)
@@
$\bar{\gamma}_{2}$ $:=\tfrac{\epsilon/(3L)}{L^{2}2\left(1+2\mathbf{k}_{\text{prd}}\right)\theta_{\rho_{\textsc{av}}}}$ (34a)
$\bar{\gamma}_{3}$ $:=\tfrac{\epsilon/(3L)}{2L(1+2\mathbf{k}_{\text{prd}})\exp(L\theta_{\rho_{\textsc{av}}})}$ (34b)
$\bar{\gamma}_{4}$ $:=\tfrac{\epsilon/(3L)}{2L\mathbf{k}_{\text{prd}}}$ (34c)
$\bar{\gamma}_{1}$ $:=\min\{\bar{\gamma}_{2},\bar{\gamma}_{3},\bar{\gamma}_{4}\}.$ (34d)
Subsequently, we pick $\gamma\in(0,\bar{\gamma}_{1})$ such that $\bar{k}:=\frac{\theta_{\rho_{\textsc{av}}}}{\gamma}\in\mathbb{N}$. This can be done without loss of generality since $\theta_{\rho_{\textsc{av}}}$ is a design parameter. Then, the definition of $\bar{\gamma}_{1}$ (cf. (34d)) and the inequality in (32) lead to the bound
$\left\|\chi^{k}-\chi^{k}_{\textsc{av}}\right\|\leq\frac{\epsilon}{L},$ (35)
@@
$\stackrel{{\scriptstyle(b)}}{{\leq}}c_{0}+\epsilon$
$\stackrel{{\scriptstyle(c)}}{{\leq}}c_{1},$ (36)
where in $(a)$ we use the fact that $\chi^{k}_{\textsc{av}}\in\Omega_{c_{0}}$ for all $k\geq 0$ (see (4) by Assumption 3) and the bound (27f), in $(b)$ we use the bound (35), while in $(c)$ we use the fact that $\epsilon\leq c_{1}-c_{0}$. Therefore, the bound (36) allows us to claim that
${\chi}^{k}\in\Omega_{c_{1}},$
for all $k\in\{0,\dots,\bar{k}\}$, i.e., we have verified that the bounds (27) can be used into the interval $\{0,\dots,\bar{k}\}$. Moreover, the exponential law (4) and the expression of $\theta_{\rho_{\textsc{av}}}$ (cf. (33)) ensure that it holds
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\rho_{\textsc{av}}+L\left\|\chi^{\bar{k}}-\chi^{\bar{k}}_{\textsc{av}}\right\|$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\rho_{\textsc{av}}+\epsilon$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\rho,$ (38)
@@
$\left\|\chi^{k+\bar{k}}-\psi_{\textsc{av}}(\chi^{\bar{k}},k+\bar{k})\right\|\leq\frac{\epsilon}{L},$
for all $k\in\{0,\dots,\bar{k}\}$. Moreover, with the same arguments, it holds $\psi_{\textsc{av}}(\chi^{\bar{k}},2\bar{k})\in\Omega_{\rho_{\textsc{av}}}$. Then, in light of Assumption 3, we guarantee that the averaged system (3) cannot escape from the set $\Omega_{\rho}$, namely
$\chi^{k}_{\textsc{av}}\in\Omega_{\rho},$
for all $k\in\{\bar{k},\dots,2\bar{k}\}$. Thus, we get $\chi^{k}\in\Omega_{\rho}$ for all $k\in\{\bar{k},\dots,2\bar{k}\}$. The proof follows by recursively applying the above arguments for each time interval $\{j\bar{k},\dots,(j+1)\bar{k}\}$ with $j=2,3,\dots$.
@@
We note that [54, Lemma 1] provides the same results claimed in Lemma 2. The only difference is that, in the mentioned reference, the objective function is assumed to have globally Lipschitz gradients. However, since we assumed compactness of the set $\mathcal{S}\subset\mathcal{K}$ and since $J$ and its gradient $G$ are continuous and differentiable [24] over the set of stabilizing gains $\mathcal{K}$, the constant
$\beta_{\mathcal{S}}$ $:=\max_{K\in\mathcal{S}}\left\|\nabla G(K)\right\|,$
is finite, and we can use it as Lipschitz constant of $G$ over $\mathcal{S}$. With this constant at hand, we can repeat all the steps in [54, Lemma 1] to conclude the proof.
@@
for all $K\in\Omega^{K}_{c_{0}}$, where $\beta_{p}:=\beta_{e}+2\sum_{\tau=k+1}^{k+\mathbf{k}_{\text{prd}}}\|\mathrm{D}^{\tau}\|/\mathbf{k}_{\text{prd}}$. Hence, by using (41), (43), and the triangle inequality, we can write the bound
@@
$\beta_{1}$ $:=\max_{K_{\textsc{av}}\in\Omega^{K}_{\tilde{c}_{0}}}\left\|G(K_{\textsc{av}})\right\|$ (47a)
$\beta_{2}$ $:=\max_{K_{\textsc{av}}\in\Omega^{K}_{\tilde{c}_{0}}}\left\|\nabla G(K_{\textsc{av}})\right\|.$ (47b)
Since also $\Omega^{K}_{\tilde{c}_{0}}$ is compact [24, Corollary 3.7.1] and recalling the continuity and differentiability of $J$ and $G$ [24], $\beta_{1}$ and $\beta_{2}$ are finite. Next, we will use them to show that ×$\mathbb{R}\times\Omega^{K}_{c_{0}}$ is (forward) invariant for (16). To this end, assume that $K_{\textsc{av}}\in\Omega^{K}_{c_{0}}$ and let us prove such an invariance using an induction argument. The increment $\Delta V_{K_{\textsc{av}}}(K_{\textsc{av}})$ of $V_{K_{\textsc{av}}}(K_{\textsc{av}}^{k})$ along trajectories of (16b) is given by

### v3

@@ V-B Averaged System Analysis
$\sum_{\tau=k+1}^{k+\mathbf{k}_{\text{prd}}}(J_{T}(K+\delta\mathrm{D}^{\tau})-\mathrm{z})\mathrm{D}^{\tau}$
@@
$\leq(1-\gamma a)^{k}V_{\lambda}(\mathrm{z}_{\textsc{av}}^{0}-J_{\textsc{av}}^{\delta}(K_{\textsc{av}}^{0}),K_{\textsc{av}}^{0})+\rho_{\textsc{av}},$ (20)
for all $k\in\mathbb{N}$. $\blacksquare$
@@ V-C Proof of Theorem 2
The proof relies on the application of Theorem 1 (cf. Section II) to system (11). Then, in order to apply Theorem 1, we need to (i) choose the design parameters $c_{1},\rho>0$ bounding the initial and final values of $V_{\lambda}$, respectively, and (ii) satisfy the conditions required by Assumptions 12, and 3. By [51, Lemma 3.8], we recall that there exists $\psi>0$ such that
$\psi\left\|K-K^{\star}\right\|^{2}\leq J(K)-J(K^{\star}),$ (21)
for all $K\in\mathcal{K}$. Therefore, by looking at the statement of Theorem 1 and given the desired final radius $r$, we set $\rho\in(0,\sqrt{r\psi}]$. In order to set the initial radius, we need to find a bound for $\delta$ such that $K^{0}+\delta\mathrm{D}^{k}$ is stabilizing for all $k\in\mathbb{N}$. To this end, we note that $K^{0}\in\mathcal{K}$, $\mathcal{K}$ is open [49, Lemma IV.3], and $\mathrm{D}^{k}$ is bounded for all $k\in\mathbb{N}$. Hence, there exists $\bar{\delta}_{0}>0$ such that $K^{0}+\delta\mathrm{D}^{k}\in\mathcal{K}$ for all $\delta\in[0,\bar{\delta}_{0}]$ and $k\in\{1,\ldots,\mathbf{k}_{\text{prd}}\}$. Now, we arbitrarily choose $c_{1}>\max_{\delta\in[0,\bar{\delta}_{0}]}V_{1}(\mathrm{z}^{0}-J_{\textsc{av}}^{\delta}(K^{0}),K^{0})$ and, thus, we note that $c_{1}\geq\max_{\delta\in[0,\bar{\delta}]}V_{\lambda}(\mathrm{z}^{0}-J_{\textsc{av}}^{\delta}(K^{0}),K^{0})$ for all $\lambda\geq 1$ (see the definition of $V_{\lambda}$ in (19)). Once the initial and final radius $c_{1}$ and $\rho$ have been chosen, let us check Assumptions 12, and 3. First, Assumption 1 is trivially satisfied because the dither signals are $\mathbf{k}_{\text{prd}}$-periodic (cf. Assumption 6). Second, we remark that (11) and its corresponding averaged system (17) are continuous over the set $\{(\mathrm{z},K)\in\mathbb{R}\times\mathcal{K}\mid K+\delta\mathrm{D}^{k}\in\mathcal{K}\text{ for all }k\in\mathbb{N}\}$, as required by Assumption 2. For this reason, let us choose $\delta$ such that the level set $\Omega_{c_{1}}:=\{(\mathrm{z},K)\in\mathbb{R}\times\mathcal{K}\mid V_{1}(\mathrm{z}-J_{\textsc{av}}^{\delta}(K),K)\leq c_{1}\}$ of $V_{1}$ (i.e, the function $V_{\lambda}$ with $\lambda=1$, see (19)) is contained into $\{(\mathrm{z},K)\in\mathbb{R}\times\mathcal{K}\mid K+\delta\mathrm{D}^{k}\in\mathcal{K}\text{ for all }k\in\mathbb{N}\}$. To this end, by looking at the definition of $V_{\lambda}$ (cf. (19)), we note that
$V_{1}(\mathrm{z}-J_{\textsc{av}}^{\delta}(K),K)\leq c_{1}\implies J(K)-J(K^{\star})\leq c_{1},$
@@
$(\mathrm{z},K)\in\Omega_{c_{1}}\implies K\in\mathcal{K}.$ (22)
Moreover, we recall that $\mathcal{K}$ is open [49, Lemma IV.3] and $\mathrm{D}^{k}$ is bounded for all $k\in\mathbb{N}$. Then, we guarantee the existence of $\bar{\delta}_{2}>0$ such that, for all $\delta\in(0,\min\{\bar{\delta}_{0},\bar{\delta}_{2}\})$, it holds $K+\delta\mathrm{D}^{k}\in\mathcal{K}$ for all $K$ satisfying $J(K)-J(K^{\star})\leq c_{1}$ and $k\in\mathbb{N}$. With these results at hand, Lemma 3 ensures the existence of $\bar{\gamma}_{1},\bar{\delta}_{1}>0$, $\bar{\lambda}\geq 1$, and $\bar{T}\in\mathbb{N}$ such that, by setting $\gamma\in(0,\bar{\gamma}_{1})$, $\delta\in(0,\bar{\delta})$ with $\bar{\delta}:=\min\{\bar{\delta}_{0},\bar{\delta}_{1},\bar{\delta}_{2}\}$, $T\geq\bar{T}$ and $\lambda\geq\bar{\lambda}$, $V_{\lambda}$ achieves the convergence properties (20) along the trajectories of the averaged system (17) and, thus, Assumption 3 is satisfied. Hence, we are entitled to apply Theorem 1 which, for all $(\mathrm{z}^{0},K^{0})\in\Omega_{c_{0}}$, ensures the existence of $\bar{\gamma}>0$ such that, for all $\gamma\in(0,\bar{\gamma})$, the trajectories of (11) satisfy
$V_{\lambda}(\mathrm{z}^{k}-J(K^{k}),K^{k})$ $\leq c_{1}$ (23a)
$V_{\lambda}(\mathrm{z}^{k}-J(K^{k}),K^{k})$ $\leq(1-\gamma a)^{k}V_{\lambda}(\mathrm{z}^{0}-J(K^{0}),K^{0})+\rho,$ (23b)
for all $k\in\mathbb{N}$. The proof of (13a) follows by combining (22) and (23a), while the proof of (13b) follows by combining (21), $J(K)-J(K^{\star})\leq V_{\lambda}(\mathrm{z}-J(K),K)$, (23b), the choice of $\rho$, and by setting $a_{0}:=\sqrt{V_{\lambda}(\mathrm{z}^{0}-J(K^{0}),K^{0})/\psi}$.
@@ VI Numerical Simulations: Control of a Doubly Fed Induction Motor
In this section, we numerically test the effectiveness of EXP-LQR and compare it with the Model-Free Policy Gradient (MFPG) method proposed in [23]. To this end, we consider a forward Euler discretization of the continuous-time linear model provided by [52] for a Doubly Fed Induction Motor (DFIM) operating at constant speed. Namely, we consider the discrete-time linear system
@@
where $\Delta t=10^{-2}$ is the adopted sampling period, while $x,u\in\mathbb{R}^{4}$ are the state and input variables and are defined as
$x:=\begin{bmatrix}i_{1u}&i_{1v}&i_{2u}&i_{2v}\end{bmatrix}^{\top},\hskip 2.84544ptu:=\begin{bmatrix}u_{1u}&u_{1v}&u_{2u}&u_{2v}\end{bmatrix}^{\top}\!\!,$
@@
$L_{1}$ $0.02645$ $R_{1}$ [$\Omega$] $0.036$
$L_{2}$ $0.0264$ $R_{2}$ [$\Omega$] $0.038$
$L_{m}$ $0.0257$ $\omega_{0}$ [rad/s] $2\pi 70.8$
@@
For the cost matrices $Q\in\mathbb{R}^{4\times 4}$ and $R\in\mathbb{R}^{4\times 4}$, we randomly generate them to ensure they are symmetric, with eigenvalues lying within the interval $(0,2)$ thus satisfying Assumption 4. We empirically tune the algorithm parameters as $\gamma=10^{-7}$, $\delta=10^{-2}$, and $T=20$ (see Remark 2). Namely, the algorithm needs to perform $4$ experiments or simulations given by $20$ samples per iteration to retrieve the truncated cost $J_{T}(K^{k}+\delta\mathrm{D}^{k})$ (see (9) and (10)). As for the generation of the dither matrix $\mathrm{D}^{k}\in\mathbb{R}^{4\times 4}$, we ordered the pairs $(i,j)\in\{1,\ldots,4\}\times\{1,\ldots,4\}$ with indices $p=1,\dots,16$ and chosen $\mathbf{k}_{p,\text{prd}}=19\times 2^{(1-p)/2}$ and $\phi_{p}=0$ for $p$ odd, while $\mathbf{k}_{p,\text{prd}}=\mathbf{k}_{p-1,\text{prd}}$ and $\phi_{p}=\pi/2$ for $p$ even. This choice ensures that Assumption 6 is satisfied with period $\mathbf{k}_{\text{prd}}=19$ by directly applying [53, Theorem]. As for MFPG, we adopt a tuning analogous to that of EXP-LQR. Specifically, we run MFPG using the same step size, perturbation amplitude, number of samples per experiment, number of experiments per iteration, and variables’ initialization used in EXP-LQR. For the sake of fairness, we also report the results achieved by MFPG with $\gamma=10^{-8}$. Fig. 4 shows the evolution of the relative cost error $(J(K^{k})-J(K^{\star}))/J(K^{\star})$ along the algorithms’ iterations $k$ in logarithmic scale. As predicted by Theorem 2, Fig. 4 shows that EXP-LQR asymptotically converges in a neighborhood of the optimal gain $K^{\star}$. Moreover, Fig. 4 shows that EXP-LQR outperforms MFPG in terms of both convergence speed and final accuracy.
Figure 4: Comparison between EXP-LQR and MFPG [23] in terms of the relative cost error $(J(K^{k})-J(K^{\star}))/J(K^{\star})$ along the algorithms’ iteration $k$.
In Fig. 5, we show the evolution of $\sigma_{\text{max}}(A+B(K^{k}+\delta\mathrm{D}^{k}))$ along the algorithm iterations $k$, where, given a generic square matrix $M\in\mathbb{R}^{n\times n}$, the symbol $\sigma_{\text{max}}(M)$ denotes the maximum (in absolute value) eigenvalue of $M$. In particular, Fig. 5 shows that $\sigma_{\text{max}}(A+B(K^{k}+\delta\mathrm{D}^{k}))$ never reaches the unitary value. Namely, as predicted by Theorem 2 (cf. (13a)), we always test the system through a stabilizing state-feedback controller $u_{t}=(K^{k}+\delta\mathrm{D}^{k})x_{t}$.
@@
Finally, in Fig. 6, we show the evolution of the norm of the state trajectory $\left\|x_{t}\right\|$ of system (24) in four simulations (each composed of $T=20$ samples) performed at different algorithm iterations $k$ to retrieve the truncated cost $J_{T}(K^{k}+\delta\mathrm{D}^{k})$. In particular, Fig. 6 shows that the trajectories of system (24) (controlled with $u_{t}=(K^{k}+\delta\mathrm{D}^{k})x_{t}$) exponentially converge to the origin quicker and quicker as the iteration index $k$ increases since we are iteratively reducing the absolute values of the eigenvalues of the gain closed-loop matrix $(A+BK^{k})$ (see also Fig. 5).
@@ VII Conclusions
We proposed EXP-LQR, i.e., a novel data-driven method able to iteratively find the state feedback gain matrix solving a Linear Quadratic Regulator problem. EXP-LQR does not need the direct knowledge of the system and cost matrices. Indeed, given an oracle able to provide a finite-time truncation of the LQR cost, our method refines its estimate according to a mechanism based on extremum-seeking. We analyzed the resulting time-varying algorithm by exploiting system theory tools based on Lyapunov stability and averaging theory. Specifically, we guaranteed that our algorithm exponentially converges to an arbitrarily small ball containing the optimal gain matrix. We tested the proposed solution with numerical simulations involving the control of an inductance motor.
@@
$\left\|f(\chi,k)\right\|$ $\leq L,\quad\left\|f_{\textsc{av}}(\chi)\right\|\leq L$ (27a)
$\left\|\dfrac{\partial f(\chi,k)}{\partial\chi}\right\|$ $\leq L,\quad\left\|\nabla f_{\textsc{av}}(\chi)\right\|\leq L$ (27b)
$\left\|\nabla V(\chi)\right\|$ $\leq L,$ (27c)
@@
$\bar{\gamma}_{2}$ $:=\tfrac{\epsilon/(3L)}{L^{2}2\left(1+2\mathbf{k}_{\text{prd}}\right)\theta_{\rho_{\textsc{av}}}},\quad\bar{\gamma}_{3}:=\tfrac{\epsilon/(3L)}{2L(1+2\mathbf{k}_{\text{prd}})\exp(L\theta_{\rho_{\textsc{av}}})}$ (35a)
[…]
Subsequently, we pick $\gamma\in(0,\bar{\gamma})$ such that $\bar{k}:=\frac{\theta_{\rho_{\textsc{av}}}}{\gamma}\in\mathbb{N}$. This can be done without loss of generality since $\theta_{\rho_{\textsc{av}}}$ is a design parameter. Then, the definition of $\bar{\gamma}$ (cf. (35b)) and the inequality in (33) lead to the bound
$\left\|\chi^{k}-\chi^{k}_{\textsc{av}}\right\|\leq\epsilon/L,$ (36)
@@
$\stackrel{{\scriptstyle(b)}}{{\leq}}c_{0}+\epsilon\stackrel{{\scriptstyle(c)}}{{\leq}}c_{1},$ (37)
where in $(a)$ we use the fact that $\chi^{k}_{\textsc{av}}\in\Omega_{c_{0}}$ for all $k\geq 0$ (see (4) by Assumption 3) and the bound (28f), in $(b)$ we use the bound (36), while in $(c)$ we use the fact that $\epsilon\leq c_{1}-c_{0}$. Therefore, the bound (37) allows us to claim that ${\chi}^{k}\in\Omega_{c_{1}}$ for all $k\in\{0,\dots,\bar{k}\}$, i.e., we have verified that the bounds (28) can be used in the interval $\{0,\dots,\bar{k}\}$. Further, the exponential law (4) and the expression of $\theta_{\rho_{\textsc{av}}}$ (cf. (34)) ensure that it holds
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\rho_{\textsc{av}}+L\left\|\chi^{\bar{k}}-\chi^{\bar{k}}_{\textsc{av}}\right\|\stackrel{{\scriptstyle(b)}}{{\leq}}\rho_{\textsc{av}}+\epsilon\stackrel{{\scriptstyle(c)}}{{\leq}}\rho,$ (39)
@@
$\left\|\chi^{k+\bar{k}}-\psi_{\textsc{av}}(\chi^{\bar{k}},k+\bar{k})\right\|\leq\epsilon/L,$
for all $k\in\{0,\dots,\bar{k}\}$. Moreover, with the same arguments, it holds $\psi_{\textsc{av}}(\chi^{\bar{k}},2\bar{k})\in\Omega_{\rho_{\textsc{av}}}$. Then, in light of Assumption 3, we guarantee that the averaged system (3) cannot escape from the set $\Omega_{\rho}$, namely, for all $k\in\{\bar{k},\dots,2\bar{k}\}$, it holds
$\chi^{k}_{\textsc{av}}\in\Omega_{\rho}.$
Thus, we get $\chi^{k}\in\Omega_{\rho}$ for all $k\in\{\bar{k},\dots,2\bar{k}\}$. The proof follows by recursively applying the above arguments for each time interval $\{j\bar{k},\dots,(j+1)\bar{k}\}$ with $j=2,3,\dots$.
@@
We note that [56, Lemma 1] provides the same results claimed in Lemma 2. The only difference is that, in the mentioned reference, the objective function is assumed to be globally $\mathcal{C}^{3}$. However, since we assumed compactness of the set $\mathcal{S}\subset\mathcal{K}$ and since $J$ and its gradient $G$ are continuously differentiable [24] over the set of stabilizing gains $\mathcal{K}$, we can repeat all the steps in [56, Lemma 1] to get the proof over $\mathcal{S}$.
@@
for all $K\in\Omega^{K}_{c_{0}}$, where $\beta_{p}:=\beta_{e}+2\sum_{\tau=k+1}^{k+\mathbf{k}_{\text{prd}}}\|\mathrm{D}^{\tau}\|/\mathbf{k}_{\text{prd}}$. Hence, by using (42), (44), and the triangle inequality, we get
@@
$\beta_{1}$ $:=\max_{K_{\textsc{av}}\in\Omega^{K}_{\tilde{c}_{0}}}\left\|G(K_{\textsc{av}})\right\|,\hskip 4.26773pt\beta_{2}:=\max_{K_{\textsc{av}}\in\Omega^{K}_{\tilde{c}_{0}}}\left\|\nabla G(K_{\textsc{av}})\right\|\!.$ (48)
Since also $\Omega^{K}_{\tilde{c}_{0}}$ is compact [24, Corollary 3.7.1] and recalling the continuity and differentiability of $J$ and $G$ [24], $\beta_{1}$ and $\beta_{2}$ are finite. Next, we will use them to show that ×$\mathbb{R}\times\Omega^{K}_{c_{0}}$ is forward-invariant for (17). To this end, assume that $K_{\textsc{av}}\in\Omega^{K}_{c_{0}}$ and let us use an induction argument. The increment $\Delta V_{K_{\textsc{av}}}(K_{\textsc{av}})$ of $V_{K_{\textsc{av}}}(K_{\textsc{av}}^{k})$ along trajectories of (17b) is given by